# Weaker Assumptions for Asymmetric Trust

### v1

@@ 5 Depth
In the symmetric world, processes are classified as correct or faulty. This suffices to express the properties of algorithms and capture all potential states of a process. In the asymmetric world this is not enough, a process might be correct but have incorrect trust assumptions. Alpos et al. [2] address this problem by introducing the notions of naive, wise, and guilds.
Existing protocols for reliable broadcast and consensus in the Alpos et al. [2] model require a guild to ensure safety and liveness. The results of Section 4, coupled with the inherent rigidity of having this kind of assumptions, motivate the search for new asymmetric algorithms that forego such assumptions. In order to do this, we introduce a new way to describe the quality of processes, based on how much they can rely on other processes during an execution. We denote this concept as the depth of a process. We then show how it can be used to characterize asymmetric problems in a more fine-grained manner. Definition 5.1 introduces this notion. This approach allows to generalize the concept of wise process or guild member and introduces more expressivity to the system.
Definition 5.1 (Depth of a process).
@@
Note that a process with maximal depth $d$ also has depth $d^{\prime}$ for all $0\leq d^{\prime}\leq d$. We will focus on the maximal depth of a process. In the following sections, any reference to the depth of a process should be understood as its maximal depth. Following the terminology of Alpos et al. [2], naive processes have depth 0, wise processes have depth at least 1, and processes in a guild have infinite depth.
@@ 6 Reliable Broadcast
Section 4 shows that if an execution includes a guild, we can convert asymmetric trust assumptions into symmetric ones. This means that if a problem can only be solved in executions with a guild, the interest in using asymmetric trust to solve it decreases. So far, the only known protocols for reliable broadcast and consensus in the Alpos et al. [2] model require the existence of a guild. This motivates the search for new algorithms that work without relying on guild assumptions for these problems. We start this search with the Byzantine reliable broadcast problem. Our approach weakens the trust assumptions needed and eliminates the need for a guild. Definition 6.1 presents a way to characterize reliable broadcast based on the depth of processes. Its properties are specific to processes that have at least a certain maximal depth.
Definition 6.1 (Depth-characterized asymmetric Byzantine reliable broadcast).
A protocol for asymmetric Byzantine reliable broadcast with sender $p_{s}$ and depth $d$, shortened as RB[$d$], defined through the events dar-broadcast(m) and dar-deliver(m) , satisfies the following properties:
@@
The reliable broadcast proposed by Alpos et al. [2] is closely related to RB[$\infty$], although not as general. RB[$\infty$] guarantees a solution for all processes with depth $\infty$, whereas the protocol by Alpos et al. [2] only guarantees a solution for those processes with depth $\infty$ that also belong to the maximal guild. That is, processes with depth $\infty$ that are members of a guild, but not of the maximal guild, are not guaranteed to receive a solution.
@@
A natural question arises, which is the minimum depth needed to implement the different primitives. A low depth not only corresponds to weaker asymmetric assumptions, but it also reflects, to a certain level, the communication complexity of an algorithm.
@@
Lemmas 6.1 and 6.2 guide the search for an algorithm. Lemma 6.1 shows that if an algorithm solves RB[$s$] then it also solves RB[$s^{\prime}$] for all $s^{\prime}\geq s$. This simplifies the search for a solution by reducing it to finding an algorithm that works for the minimal value of $s$. Lemma 6.2 shows that there are no algorithms that can solve RB[1], therefore, we must search for a protocol that solves the problem for processes with a depth of at least 2.
Lemma 6.1.
Proof.
Lemma 6.2.
There is no quorum-based algorithm that solves RB[1]
Proof.
Firstly consider an execution $E_{0}$ in which $p_{5}$ dar-broadcasts value 0, all processes behave according to the algorithm. Processes $p_{1}$ and $p_{2}$ will deliver value 0. From Definition 3.5 we know that there is an equivalent execution $E$ where $p_{1}$ and $p_{2}$ deliver the same value and only receive messages from the processes in their quorum $\{4,5,6\}$ and $p_{5}$.
Now consider an execution $E^{\prime}$ in which processes $p_{5}$ and $p_{6}$ behave towards $p_{4}$ like in $E$. Note that $p_{4}$ will behave like in $E$ since at every step it will receive the same messages from the same processes as it did in $E$. Let $p_{5}$ and $p_{6}$ also behave towards $p_{2}$ like in $E$ and send no messages towards all other processes. Process $p_{3}$ does not begin the execution since it never receives any message from the initiator (Byzantine) nor any member of one of its quorums (processes 3, 5 or 6). Process $p_{2}$ will deliver value 0 as it receives the same messages from the same processes as in $E$. On the contrary, process $p_{1}$ will not deliver any value since it will never receive messages from all processes in one of its quorums, which would contradict Definition 3.5. This comes from the fact that $p_{3}$ will never send any message and that $p_{5}$ and $p_{6}$ are Byzantine and will not send it any message. Therefore, $p_{2}$ delivers a value but $p_{1}$ does not, which breaks the totality property of RB[1]. ∎
@@
Algorithm 1 presents a solution for RB[3]. Every process waits to receive a quorum of Readyafterecho messages associated to the same message $m$ and round $r$ before delivering $m$. A process with depth 3 will only receive such a quorum if there exists a set of processes $Q$ (which form a quorum for a wise process) that can attest that the sender indeed sent the value $m$. Since there will be an attesting quorum for every process that delivers, and since quorums for wise processes intersect in at least one correct process, we can deduce that all processes with depth 3 that deliver a value $m$ will deliver the same value. In addition, we use a technique similar to Bracha’s kernel amplification [7] to ensure that if a valid process delivers then all valid processes will deliver. The arrays in lines 1,  1,  1, and 1 are hashmaps, so even though they are depicted as having infinite size they are actually sparsely populated. The algorithm has a latency of 3 asynchronous rounds in the good case and 5 in the bad case. Its correctness is proved in Lemma 6.3. The existence of an algorithm to solve RB[2] remains an open question.
@@
Lemma 6.3.
@@
Consistency: Suppose a process $p_{i}$ with depth 3 dar-delivered a value $m_{i}$ and another process with depth 3 $p_{j}$ dar-delivered a value $m_{j}$. By Lemma 6.6 this means that there exists a process with depth 1 $p_{i}^{\prime}$ such that it received $\textsc{Echo},{m_{i}}]$ messages from one of its quorums $Q_{i^{\prime}}\in\mathcal{Q}_{i^{\prime}}$ and another process with depth 1 $p_{j}^{\prime}$ that received $\textsc{Echo},{m_{j}}]$ messages from one of its quorums $Q_{j^{\prime}}\in\mathcal{Q}_{j^{\prime}}$. Since $Q_{i}^{\prime}\cap Q_{j}^{\prime}\nsubseteq F$ there is at least one correct process that sent $\textsc{Echo},{m_{i}}]$ and $\textsc{Echo},{m_{j}}]$ messages. But since a correct process will only send one type of echo messages, we conclude that $m_{i}=m_{j}$.
@@
Since all processes with depth 1 send a $\textsc{Readyafterready},{n,m}]$ message, all processes with depth 2 will receive such messages from one of their quorums, prompting them to send a
$\textsc{Readyafterecho},{n+1,m}]$ message. Since all processes with depth 2 will send such messages all processes with depth 3 will receive $\textsc{Readyafterecho},{n+1,m}]$ from one of their quorums and proceed to dar-deliver the value $m$.
Corollary 6.1.
@@
Lemma 6.4.
In any execution with at least one process $p_{3}$ with depth 3 and for any $n>1$, if a process $p_{i}$ with depth 1 sends a $\textsc{Readyafterecho},{n,m}]$ message then there exists another process $p_{j}$ with depth 1 that sent a
$\textsc{Readyafterecho},{n-1,m}]$ message.
Proof.
According to the code of Algorithm 1, $p_{i}$ sends a
$\textsc{Readyafterecho},{n,m}]$ message after receiving
$\textsc{Readyafterready},{n-1,m}]$ messages from one of its quorums $Q_{i}\in\mathcal{Q}_{i}$. Process $p_{3}$ has a quorum $Q_{2}\in\mathcal{Q}_{3}$ such that all its members have depth at least 2. By the quorum consistency property there is at least one process $p_{2}\in Q_{i}\cap Q_{2}$. Therefore, $p_{2}$ sent a $\textsc{Readyafterready},{n-1,m}]$ message.
According to Algorithm 1, $p_{2}$ does this after receiving
$\textsc{Readyafterecho},{n-1,m}]$ messages from a kernel $K_{2}\in\mathcal{K}_{2}$. Since $p_{2}$ has depth 2, it has at least one quorum where all its members have depth 1, and therefore each of its kernels contains at least one process with depth 1. Let $p_{j}$ be such a process in $K_{2}$. Since $p_{j}$ has depth 1 and it sent a $\textsc{Readyafterecho},{n-1,m}]$ message we have proven the lemma. ∎
Lemma 6.5.
Proof.
According to Algorithm 1, a correct process only sends a
$\textsc{Readyafterecho},{1,m}]$ message after receiving $\textsc{Echo},{m}]$ messages from one of its quorums $Q_{i}\in\mathcal{Q}_{i}$. This proves the lemma. ∎
Lemma 6.6.
Proof.
Since a process with depth 2 also has depth 1, we can apply Lemma 6.4 to obtain that there is another process $p_{j}^{\prime}$ with depth 1 that sent a
$\textsc{Readyafterecho},{n-1,m}]$ message. By repeatedly applying Lemma 6.4 we reach a a process $p_{j}$ with depth 1 that sent a $\textsc{Readyafterecho},{1,m}]$ message. We now apply Lemma 6.5 to obtain that $\exists Q_{j}\in\mathcal{Q}_{j}$ such that all processes in $Q_{j}$ sent a $\textsc{Echo},{m}]$ message. This proves the lemma. ∎
In Algorithm 1 and Definition 6.1, we assume that all correct processes start the execution. This allows us to focus solely on characterizing the depth of the processes that complete the execution. However, in more complex scenarios, such as when protocols are composed sequentially, it is important to characterize the depth of the processes that start the execution.
@@
In the case of reliable broadcast, we treat it as a standalone algorithm and thus omit this distinction. However, in later sections, we will define protocols using two depth parameters: one for the processes that begin execution, and one for those that finish.
@@ 7 Consensus
We adapt the consensus problem to the depth paradigm and propose an algorithm that solves it for all processes with a minimum depth of at least 9. We follow an approach similar to the one proposed by Abraham et al. [1], using a round-based algorithm composed of a binding crusader agreement and a common coin. For this purpose, we introduce depth-characterized versions of these primitives in Sections 7.1 and 7.2, which we use as subprotocols inside the main consensus algorithm.
It is at this point that the need to characterize the depth of starting and finishing processes arises. The consensus protocol first calls the binding crusader agreement, and then the common coin. Only processes that finish the first call will be able to make the call to the second protocol. Because of this, we need to take into account that not all correct processes can start the execution. We introduce two parameters to characterize the depth of protocols: $d^{\prime}$ to refer to the depth of processes that start the execution, and $d$ to refer to the depth of processes that finish the execution. We will say that an algorithm works if, in case all processes with depth $d^{\prime}$ start the execution, all processes with depth $d$ satisfy the required properties. In case we assume that all correct processes start (i.e., $d^{\prime}=0$), as is the case for standalone protocols (like consensus and reliable broadcast), we will ignore the $d^{\prime}$ parameter.
@@ 7.1 Common Coin
The existing asymmetric common coin protocol proposed by Alpos et al. [2] has issues that motivate the development of a new protocol. It is not a purely asymmetric protocol, as each process first constructs the tolerated system, which is a symmetric fail-prone system, and executes the protocol based upon it. This means that the protocol is actually symmetric and is not suitable for asymmetric applications. We address this by first redefining the problem to fit the depth paradigm, shown in Definition 7.1, and propose a modified version of the scheme introduced by Alpos et al. [2] that solves the problem for all processes with depth at least 7, given that all processes with depth 6 start the execution. The parameter for all processes starting the protocol is needed since it will be used as part of a pipeline of protocols, which means that only the processes that finish the preceding protocol are guaranteed to start executing the common coin.
Definition 7.1 (Depth-characterized asymmetric common coin).
A protocol for common coin, for processes with depth $d^{\prime}$ and $d$, defined by the events release-coin(round) and output-coin(c, round), shortened as CC[d,d][$^{\prime},d$], satisfies the following properties:
@@
If a process with depth $d$ executes output-coin(c, round) and another process with depth $d$ executes output-coin(c, round), then $\textit{c}=\textit{c}^{\prime}$.
@@
Algorithm 2 provides a solution to the common coin taking $d=7$ and $d^{\prime}=6$. Note that these values are not the smallest depths that allow the algorithm to solve the problem, but are the depths we will need within the context of the consensus algorithm. In fact, the same algorithm allows to solve CC[0, 1]. The common coin scheme follows the approach of Rabin [23] and assumes that coins are predistributed by a trusted dealer. The scheme uses Benaloh-Leichter [6] secret sharing, such that the coin is additively shared within every quorum. The dealer shares one coin for every possible round of the protocol and when a process receives a share it can verify that it was indeed created by the dealer. This can be done using standard digital signatures but is outside the scope of this paper. Alpos et al. [2] follow the same outline in their scheme.
@@
for all $p_{j}\in\mathcal{P}$ do
send message $\textsc{share},{s_{iQ},Q}]$ to $p_{j}$
@@
Lemma 7.2.
Algorithm 2 implements a depth-characterized common coin for $d=7$ and $d^{\prime}=6$.
@@
Termination: Consider a process $p_{i}$ with depth 7. There exists a quorum $Q\in\mathcal{Q}_{i}$ such that all its members have depth at least 6. Since we assume that they all start the protocol, they will release the coin and send their share to all the processes in $\mathcal{P}$. Therefore, $p_{i}$ will receive all the coin shares for at least the quorum $Q$ and will be able to output the coin value.
Unpredictability: Assume a process $p_{7}$ with depth 7 outputs the coin. This implies the existence of a set $Q\in\mathcal{Q}_{7}$ such that each member of $Q$ sent a share message. Since $p_{7}$ has a quorum $Q_{7}\in\mathcal{Q}_{7}$ such that $\forall p\in Q_{7},\textsl{depth(}p\textsl{)}\geq 6$ and $Q\cap Q_{7}\neq\emptyset$, it follows that at least one process with depth 6 has released the coin.
@@
7.2 Binding Crusader Agreememt

### v2

@@ 5 Depth
In the symmetric world, processes are classified as correct or faulty. This suffices to express the properties of algorithms and capture all potential states of a process. In the asymmetric world this is not enough, a process might be correct but have incorrect trust assumptions. Damgård et al. [DBLP:conf/asiacrypt/DamgardDFN07] and Alpos et al. [DBLP:journals/dc/AlposCTZ24] address this problem by introducing the notions of naive, wise, and guilds.
Existing protocols for reliable broadcast and consensus in the Alpos et al. [DBLP:journals/dc/AlposCTZ24] model require a guild to ensure safety and liveness. The results of Section 4, coupled with the inherent rigidity of having this kind of assumptions, motivate the search for new asymmetric algorithms that forego such assumptions. In order to do this, we introduce a new way to describe the quality of processes, based on how much they can rely on other processes during an execution. We denote this concept as the depth of a process. We then show how it can be used to characterize asymmetric problems in a more fine-grained manner. Definition 7 introduces this notion. This approach allows to generalize the concept of wise process or guild member and introduces more expressivity to the system.
Definition 7 (Depth of a process).
@@
Note that a process with maximal depth $d$ also has depth $d^{\prime}$ for all $0\leq d^{\prime}\leq d$. We will focus on the maximal depth of a process. In the following sections, any reference to the depth of a process should be understood as its maximal depth. Following the terminology of Damgård et al. [DBLP:conf/asiacrypt/DamgardDFN07], naive processes have depth 0, wise processes have depth at least 1, and processes in a guild have infinite depth.
@@ 6 Reliable Broadcast
Section 4 shows that if an execution includes a guild, we can convert asymmetric trust assumptions into symmetric ones. This means that if a problem can only be solved in executions with a guild, the interest in using asymmetric trust to solve it decreases. So far, the only known protocols for reliable broadcast and consensus in the Alpos et al. [DBLP:journals/dc/AlposCTZ24] model require the existence of a guild. This motivates the search for new algorithms that work without relying on guild assumptions for these problems. We start this search with the Byzantine reliable broadcast problem. Our approach weakens the trust assumptions needed and eliminates the need for a guild. Definition 8 presents a way to characterize reliable broadcast based on the depth of processes. Its properties are specific to processes that have at least a certain maximal depth.
Definition 8 (Depth-characterized asymmetric Byzantine reliable broadcast).
A protocol for asymmetric Byzantine reliable broadcast with sender $p_{s}$ and depth $d$, shortened as RB[$d$], defined through the events dar-broadcast(m) and dar-deliver(m) satisfies the following properties:
@@
The reliable broadcast proposed by Alpos et al. [DBLP:journals/dc/AlposCTZ24] is closely related to RB[$\infty$], although not as general. RB[$\infty$] guarantees a solution for all processes with depth $\infty$, whereas the protocol by Alpos et al. [DBLP:journals/dc/AlposCTZ24] only guarantees a solution for those processes with depth $\infty$ that also belong to the maximal guild. That is, processes with depth $\infty$ that are members of a guild, but not of the maximal guild, are not guaranteed to receive a solution.
@@
A natural question arises, which is the minimum depth needed to implement the different primitives. A low depth not only corresponds to weaker asymmetric assumptions, but also reflects, to a certain level, the communication complexity of an algorithm.
@@
If an algorithm solves RB[$s$] then it also solves RB[$s^{\prime}$] for all $s^{\prime}\geq s$. This simplifies the search for a solution by reducing it to finding an algorithm that works for the minimal value of $s$. Additionally, there are no algorithms that can solve RB[1], therefore, we must search for a protocol that solves the problem for processes with a depth of at least 2. These results are proven in Lemmas 6 and 7.
@@
Algorithm 1 presents a solution for RB[3]. Every process waits to receive a quorum of Readyafterecho messages associated to the same message $m$ and round $r$ before delivering $m$. A process with depth 3 will only receive such a quorum if there exists a set of processes $Q$ (which form a quorum for a wise process) that can attest that the sender indeed sent the value $m$. Since there will be an attesting quorum for every process that delivers, and since quorums for wise processes intersect in at least one correct process, we can deduce that all processes with depth 3 that deliver a value $m$ will deliver the same value. In addition, we use a technique similar to Bracha’s kernel amplification [DBLP:journals/jacm/BrachaT85] to ensure that if a valid process delivers then all valid processes will deliver. The arrays in lines 1,  1,  1, and 1 are hashmaps, so even though they are depicted as having infinite size they are actually sparsely populated. The algorithm has a latency of 3 asynchronous rounds in the good case and 5 in the bad case. Its correctness is proved in Theorem 2. The existence of an algorithm to solve RB[2] remains an open question.
@@
Lemma 3.
In any execution with at least one process $p_{3}$ with depth 3 and for any $n>1$, if a process $p_{i}$ with depth 1 sends a $\textsc{Readyafterecho},{n,m}]$ message then there exists another process $p_{j}$ with depth 1 that sent a $\textsc{Readyafterecho},{n-1,m}]$ message.
Proof.
According to the code of Algorithm 1, $p_{i}$ sends a $\textsc{Readyafterecho},{n,m}]$ message after receiving $\textsc{Readyafterready},{n-1,m}]$ messages from one of its quorums $Q_{i}\in\mathcal{Q}_{i}$. Process $p_{3}$ has a quorum $Q_{2}\in\mathcal{Q}_{3}$ such that all its members have depth at least 2. By the quorum consistency property there is at least one process $p_{2}\in Q_{i}\cap Q_{2}$. Therefore, $p_{2}$ sent a $\textsc{Readyafterready},{n-1,m}]$ message.
According to Algorithm 1, $p_{2}$ does this after receiving $\textsc{Readyafterecho},{n-1,m}]$ messages from a kernel $K_{2}\in\mathcal{K}_{2}$. Since $p_{2}$ has depth 2, it has at least one quorum where all its members have depth 1, and therefore each of its kernels contains at least one process with depth 1. Let $p_{j}$ be such a process in $K_{2}$. Since $p_{j}$ has depth 1 and it sent a $\textsc{Readyafterecho},{n-1,m}]$ message we have proven the lemma. ∎
Lemma 4.
Proof.
According to Algorithm 1, a correct process only sends a $\textsc{Readyafterecho},{1,m}]$ message after receiving $\textsc{Echo},{m}]$ messages from one of its quorums $Q_{i}\in\mathcal{Q}_{i}$. This proves the lemma. ∎
Lemma 5.
Proof.
Since a process with depth 2 also has depth 1, we can apply Lemma 3 to obtain that there is another process $p_{j}^{\prime}$ with depth 1 that sent a $\textsc{Readyafterecho},{n-1,m}]$ message. By repeatedly applying Lemma 3 we reach a a process $p_{j}$ with depth 1 that sent a $\textsc{Readyafterecho},{1,m}]$ message. We now apply Lemma 4 to obtain that $\exists Q_{j}\in\mathcal{Q}_{j}$ such that all processes in $Q_{j}$ sent a $\textsc{Echo},{m}]$ message. This proves the lemma. ∎
Lemma 6.
Proof.
Lemma 7.
There is no quorum-based algorithm that solves RB[1]
Proof.
Firstly consider an execution $E_{0}$ in which $p_{5}$ dar-broadcasts value 0, all processes behave according to the algorithm. Processes $p_{1}$ and $p_{2}$ will deliver value 0. From Definition 5 we know that there is an equivalent execution $E$ where $p_{1}$ and $p_{2}$ deliver the same value and only receive messages from the processes in their quorum $\{4,5,6\}$ and $p_{5}$.
Now consider an execution $E^{\prime}$ in which processes $p_{5}$ and $p_{6}$ behave towards $p_{4}$ like in $E$. Note that $p_{4}$ will behave like in $E$ since at every step it will receive the same messages from the same processes as it did in $E$. Let $p_{5}$ and $p_{6}$ also behave towards $p_{2}$ like in $E$ and send no messages towards all other processes. Process $p_{3}$ does not begin the execution since it never receives any message from the initiator (Byzantine) nor any member of one of its quorums (processes 3, 5 or 6). Process $p_{2}$ will deliver value 0 as it receives the same messages from the same processes as in $E$. On the contrary, process $p_{1}$ will not deliver any value since it will never receive messages from all processes in one of its quorums, which would contradict Definition 5. This comes from the fact that $p_{3}$ will never send any message and that $p_{5}$ and $p_{6}$ are Byzantine and will not send it any message. Therefore, $p_{2}$ delivers a value but $p_{1}$ does not, which breaks the totality property of RB[1]. ∎
Theorem 2.
@@
Consistency: Suppose a process $p_{i}$ with depth 3 dar-delivered a value $m_{i}$ and another process with depth 3 $p_{j}$ dar-delivered a value $m_{j}$. By Lemma 5 this means that there exists a process with depth 1 $p_{i}^{\prime}$ such that it received $\textsc{Echo},{m_{i}}]$ messages from one of its quorums $Q_{i^{\prime}}\in\mathcal{Q}_{i^{\prime}}$ and another process with depth 1 $p_{j}^{\prime}$ that received $\textsc{Echo},{m_{j}}]$ messages from one of its quorums $Q_{j^{\prime}}\in\mathcal{Q}_{j^{\prime}}$. Since $Q_{i}^{\prime}\cap Q_{j}^{\prime}\nsubseteq F$ there is at least one correct process that sent $\textsc{Echo},{m_{i}}]$ and $\textsc{Echo},{m_{j}}]$ messages. But since a correct process will only send one type of echo messages, we conclude that $m_{i}=m_{j}$.
@@
Since all processes with depth 1 send a $\textsc{Readyafterready},{n,m}]$ message, all processes with depth 2 will receive such messages from one of their quorums, prompting them to send a $\textsc{Readyafterecho},{n+1,m}]$ message. Since all processes with depth 2 will send such messages all processes with depth 3 will receive $\textsc{Readyafterecho},{n+1,m}]$ from one of their quorums and proceed to dar-deliver the value $m$.
Corollary 1.
@@
In Algorithm 1 and Definition 8, we assume that all correct processes start the execution. This allows us to focus solely on characterizing the depth of the processes that complete the execution. However, in more complex scenarios, such as when protocols are composed sequentially, it is important to also characterize the depth of the processes that start the execution.
@@
In the case of reliable broadcast, we treat it as a standalone algorithm and thus omit this distinction. In later sections, we will define protocols using two depth parameters: one for the processes that begin the execution, and one for those that finish.
@@ 7 Consensus
We adapt the consensus problem to the depth paradigm and propose an algorithm that solves it for all processes with a minimum depth of at least 10. We follow an approach similar to the one proposed by Abraham et al. [DBLP:conf/podc/AbrahamBY22], who use a round-based algorithm composed of a binding crusader agreement and a common coin. However, for reasons that will be discussed in Section 7.3, we replace binding crusader agreement with a binding graded agreement protocol which allows parties to provide multiple input. For this purpose, we introduce depth-characterized versions of these primitives in Sections 7.1 and 7.2, which we use as subprotocols inside the main consensus algorithm.
It is at this point that the need to characterize the depth of starting and finishing processes arises. The consensus protocol first calls the binding graded agreement, and then the common coin. Only processes that finish the former will be able to make the call to the latter. Because of this, we need to take into account that not all correct processes can start executing a protocol. We introduce two parameters to characterize the depth of protocols: $d^{\prime}$ to refer to the depth of processes that start the execution, and $d$ to refer to the depth of processes that finish the execution. We will say that an algorithm works if, in case all processes with depth $d^{\prime}$ start the execution, all processes with depth $d$ satisfy the required properties. In case we assume that all correct processes start (i.e., $d^{\prime}=0$), as is the case for standalone protocols (like consensus and reliable broadcast), we will ignore the $d^{\prime}$ parameter.
@@ 7.1 Common Coin
The existing asymmetric common coin protocol proposed by Alpos et al. [DBLP:journals/dc/AlposCTZ24] has issues that motivate the development of a new protocol. It is not a purely asymmetric protocol, as each process first constructs the tolerated system, which is a symmetric fail-prone system, and executes the protocol based upon it. This means that the protocol is actually symmetric and is not suitable for asymmetric applications. We address this by first redefining the problem to fit the depth paradigm, shown in Definition 9, and propose a modified version of the scheme introduced by Alpos et al. [DBLP:journals/dc/AlposCTZ24], which solves the problem for all processes with depth at least $d+1$, as long as all processes with depth $d$ start the execution. The parameter for all processes starting the protocol is needed since it will be used as part of a pipeline of protocols, which means that only the processes that finish the preceding protocol are guaranteed to start executing the common coin.
Definition 9 (Depth-characterized asymmetric common coin).
A protocol for common coin, for processes with depth $d^{\prime}$ and $d$, defined by the events release-coin(round) and output-coin(c), shortened as CC[d,d][$^{\prime},d$], satisfies the following properties:
@@
If a process with depth $d$ executes output-coin(c) and another process with depth $d$ executes output-coin(c), then $\textit{c}=\textit{c}^{\prime}$.
@@
Algorithm 2 provides a solution to the common coin taking any value $d^{\prime}$ and $d=d^{\prime}+1$. In our consensus construction, we require CC[8, 9] in particular. The common coin we design is strong and it follows the approach of Rabin [DBLP:conf/focs/Rabin83] and assumes that coins are predistributed by a trusted dealer. The scheme uses Benaloh-Leichter [DBLP:conf/crypto/Leichter88] secret sharing, such that the coin is additively shared within every quorum. The dealer shares one coin for every possible round of the protocol and when a process receives a share it can verify that it was indeed created by the dealer. This can be done using standard digital signatures but is outside the scope of this paper. Alpos et al. [DBLP:journals/dc/AlposCTZ24] follow the same approach in their scheme. Theorem 3 proves the correctness of the algorithm.
@@
send message $\textsc{share},{s_{iQ},Q}]$ to $p_{j}$ to all $p_{j}$ in $\mathcal{P}$
@@
Theorem 3.
Algorithm 2 implements CC[d, d+1].
@@
Termination: Consider a process $p_{i}$ with depth $d+1$. There exists a quorum $Q\in\mathcal{Q}_{i}$ such that all its members have depth at least $d$. Since we assume that they all start the protocol, they will release the coin and send their share to all the processes in $\mathcal{P}$. Therefore, $p_{i}$ will receive all the coin shares for at least the quorum $Q$ and will be able to output the coin value.
Unpredictability: Assume a process $p_{i}$ with depth $d+1$ outputs the coin. This implies the existence of a set $Q\in\mathcal{Q}_{i}$ such that each member of $Q$ sent a share message. Since $p_{i}$ has a quorum $Q_{d}\in\mathcal{Q}_{i}$ such that $\forall p\in Q_{i},\textsl{depth(}p\textsl{)}\geq d$ and $Q\cap Q_{d}\neq\emptyset$, it follows that at least one process with depth d has released the coin.
@@
7.2 Binding Graded Agreement